# Complementarity-based complementarity

### v1

@@ I Complementarity of three MUBs
The bound (1) was known [9] and was implicitly assumed to hold for every possible choice of MUB triplet, whereas the bound (2), obtained numerically (see in the appendix), was found while experimentally checking the entropic uncertainty relations. The states that minimize the first bound are eigenstates of any of the three $\hat{A},\hat{B},\hat{C}$, e.g. $|\psi\rangle=|a_{0}\rangle$. The states that minimize the second bound, instead, have a single null component when expressed in the computational basis $|a_{i}\rangle$, e.g. a state $\sum_{n}\psi_{n}e^{i\phi_{n}}|a_{n}\rangle$ with $\{\psi_{n}\}=(0.54488,0.45067,0,0.45067,0.54488)$, $\{\phi_{n}\}=\left(0,-{2}\pi/{5},0,\pi,-{\pi}/{5}\right)$. To validate the numerical minimization, we calculated the minimal sum of the entropies on a large number of states chosen randomly [16, 17] with Haar measure (Fig. 2), confirming that no state beats either bound (1), (2). We also checked that the sum of any two entropies never beats the Maassen-Uffink bound [8] of $\log_{2}5$ (Fig. 3), confirming that the effect presented here requires at least three mutually unbiased bases, as expected.
@@
Interestingly, even though it is known that there are two inequivalent classes of MUBs also in dimension $d=4$ [14], we could not find a similar mismatch in the sum of entropies of three MUBs: for $d=4$ there are 5 MUBs connected to five observables ($\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}$), and we checked all $\binom{5}{3}=10$ combinations of triplets $\hat{X},\hat{Y},\hat{Z}$ in $\{\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}\}$. They all have the lower bound $H(\hat{X})+H(\hat{Y})+H(\hat{Z})\geqslant 3$ identified in [9]. However, while in the case $d=3$ the states achieving the lower bound are the same for every triplet, in $d=4$ the optimal states depend on the choice of the MUBs. We note that they share a common structure, i.e. they all have two non-null components.
By studying the sum of variances in $d=4$ and $d=5$ we find the same behavior of the uncertainty relation as the one found for the EURs: In $d=4$ the optimal states attain the same minimum for every choice of the bases, while in $d=5$ we find two distinct lower bounds dividing the triplets of MUBs in the same two sets as for the EURs. In $d=4$ all the uncertainty relations for every choice of the three MUBs read
@@ II Experimental verification
We experimentally tested the entropic results presented in the previous section by encoding information in the time-frequency degree of freedom of photons [18]. Namely, we consider a Hilbert space generated by broadband frequency bins and their superpositions encoded in coherent light pulses. In this encoding alphabet, the five-dimensional computational basis $|a_{i}\rangle$ (associated to observable $\hat{A}$) is defined as a set of five Gaussian-shaped frequency bins centered at different frequencies, and the MUBs are generated by superimposing the fundamental bins with different phases. We note that the chosen alphabet falls in the category of pulsed temporal modes as the superposition states overlap in both time and frequency.
@@ Appendix B Numerical minimization
$\Delta\hat{X}^{2}=\langle\psi|\hat{X}^{2}|\psi\rangle-\langle\psi|\hat{X}|\psi\rangle^{2},$ (11)

### v2

@@ I Complementarity of three MUBs
The bound (1) was known [9] and was implicitly assumed to hold for every possible choice of MUB triplet, whereas the bound (2), obtained numerically (see in the appendix), was found while experimentally checking the entropic uncertainty relations. The states that minimize the first bound are eigenstates of any of the three $\hat{A},\hat{B},\hat{C}$, e.g. $|\psi\rangle=|a_{0}\rangle$. The states that minimize the second bound, instead, have a single null component when expressed in the computational basis $|a_{i}\rangle$, e.g. a state $\sum_{n}\psi_{n}e^{i\phi_{n}}|a_{n}\rangle$ with $\{\psi_{n}\}=(0.54488,0.45067,0,0.45067,0.54488)$, $\{\phi_{n}\}=\left(0,-{2}\pi/{5},0,\pi,-{\pi}/{5}\right)$. To validate the numerical minimization, we calculated the minimal sum of the entropies on a large number of states chosen randomly [16, 17] with Haar measure (Fig. 2), confirming that no state beats either bound (1), (2). We also checked that the sum of any two entropies never beats the Maassen-Uffink bound [8] of $\log_{2}5$ (Fig. 3) and that the sum of $four$ entropies has the unique lower bound found in Ref. [9] for every choice of the MUBs, confirming that the effect presented here requires exactly three MUBs.
@@
Interestingly, even though it is known that there are two inequivalent classes of MUBs also in dimension $d=4$ [14], we could not find a similar mismatch in the sum of entropies of three MUBs: for $d=4$ there are 5 MUBs connected to five observables ($\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}$), and we checked all $\binom{5}{3}=10$ combinations of triplets $\hat{X},\hat{Y},\hat{Z}$ in $\{\hat{A},\hat{B},\hat{C},\hat{D},\hat{E}\}$. They all have the lower bound $H(\hat{X})+H(\hat{Y})+H(\hat{Z})\geqslant 3$ identified in Ref. [9]. However, while in the case $d=3$ the states achieving the lower bound are the same for every triplet, in $d=4$ the optimal states depend on the choice of the MUBs. We note that they share a common structure, i.e. they all have two non-null components.
By studying the sum of variances in $d=4$ and $d=5$ we find the same behavior of the uncertainty relation as the one found for the EURs: In $d=4$ the optimal states attain the same minimum for every choice of the bases, while in $d=5$ we find two distinct lower bounds dividing the triplets of MUBs in the same two sets as for the EURs. For $d=4$ all the uncertainty relations for every choice of the three MUBs read
@@ II Experimental verification
We experimentally tested the entropic results presented in the previous section by encoding information in photonic time-frequency modes [18]. Namely, we consider a Hilbert space generated by broadband frequency bins and their superpositions encoded in coherent light pulses. In this encoding alphabet, the five-dimensional computational basis $|a_{i}\rangle$ (associated to observable $\hat{A}$) is defined as a set of five Gaussian-shaped frequency bins centered at different frequencies, and the MUBs are generated by superimposing the fundamental bins with different phases. We note that the chosen alphabet falls in the category of pulsed temporal modes as the superposition states overlap in both time and frequency.
@@ Appendix B Numerical minimization
$\Delta X^{2}=\langle\psi|\hat{X}^{2}|\psi\rangle-\langle\psi|\hat{X}|\psi\rangle^{2},$ (11)